### v2

# State Your Intention to Steer Your Attention: An AI Assistant for Intentional Digital Living

Abstract.
When working on digital devices, people often face distractions that can lead to a decline in productivity and efficiency, as well as negative psychological and emotional impacts. To address this challenge, we introduce a novel Artificial Intelligence (AI) assistant that elicits a user’s intention, assesses whether ongoing activities are in line with that intention, and provides gentle nudges when deviations occur. The system leverages a large language model to analyze screenshots, application titles, and URLs, issuing notifications when behavior diverges from the stated goal. Its detection accuracy is refined through initial clarification dialogues and continuous user feedback. In a three-week, within-subjects field deployment with 22 participants, we compared our assistant to both a rule-based intent reminder system and a passive baseline that only logged activity. Results indicate that our AI assistant effectively supports users in maintaining focus and aligning their digital behavior with their intentions. Our source code is publicly available at https://intentassistant.github.io
AI Assistant, User Intention, Digital Self-Control, Distraction Reduction, Attention Management, Proactive AI, Screen-Context Understanding, Large Language Models
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interaction design
1. Introduction
Digital devices such as computers and smartphones have become essential tools for a wide range of tasks, spanning activities like studying, professional work, and planning travel. However, while using such devices, users are constantly assailed with tempting opportunities for distraction (Lyngs et al., 2019), both from the device itself and from their own thoughts of alternative on-device activities they could engage in. A variety of applications have been developed to help users stay on task (Roffarello and De Russis, 2023; Hiniker et al., 2016; Ko et al., 2015a; Kim et al., 2019; Collins et al., 2014; Ahmed et al., 2014). Yet, most remain limited to rule-based blockers or simple time trackers that operate within narrow boundaries. Such approaches fail to capture the context of what users intend to do versus what they are actually doing, and often impose excessive restrictions when the same application serves multiple purposes.
In our formative study with eleven active computer users, we found a clear need for more intelligent, context-aware assistance beyond rule-based approaches. Participants expressed a desire for a system that could interpret their often broad and ambiguous intentions, support them in staying aligned with their goals, and provide gentle nudges when deviations occur. They emphasized that such an assistant should be immediate in response and gentle in manner.
In this work, we introduce the Intent Assistant (INA), an AI system that uses large language models (LLMs) to help users remain focused on their intended tasks in digital environments. As shown in Figure 1, the user first enters their intention to INA as text, which is followed by a brief clarification dialogue to capture the goal more precisely. INA then continuously monitors on-screen activity, detects potential distractions, and delivers timely, gentle nudges when the user diverges from their stated intention. Over time, the system incorporates user feedback to refine its judgments in real time, enabling progressively more accurate and supportive interventions. By preserving user autonomy while gently guiding them back toward their goals, INA serves as a collaborative assistant in everyday digital life.
We evaluate the distraction detection capability of INA and analyze the effectiveness of our design choices. To this end, we introduce IntentionBench, a novel dataset of activity records in computer use, featuring diversity and realistic user workflows. To create IntentionBench, we collected diverse activity records by carrying out 50 unique task instructions across 14 applications and 32 websites. From these on-task sessions, we generated off-task samples by replacing user intentions from one session with those of another, yielding a dataset of approximately 77 hours of mixed on-task and off-task activities reflecting natural transition points (see Section 5.1 for more details). On IntentionBench, INA detects distraction from screenshot context with an accuracy of 0.878 and an F1-score of 0.845.
Finally, we conducted a three-week in-the-wild deployment with 22 participants, comparing INA against a simple reminder application and a logging only application. Participants using INA showed a significantly lower LLM-estimated off-task ratio (0.104 vs. 0.166 with simple reminder, $p<.001$), higher intention alignment ratings (4.44 vs. 4.23 with simple reminder, $p<.001$), and greater focused immersion (3.74 vs. 3.34 with simple reminder, $p=.045$; vs. 2.90 with logging only, $p=.0003$). Beyond the quantitative outcomes, weekly surveys and interviews not only reinforced our findings but also revealed additional benefits and opportunities for improvement. Participants regarded INA as enhancing intention fulfillment, strengthening awareness of digital habits, and providing supportive companionship. At the same time, they surfaced challenges such as notification burden, detection accuracy, and long-term user adaptation. We summarize our main contributions as follows:
  • Through a formative study with frequent computer users, we identify limitations of existing productivity tools and uncover users’ needs for an assistant that supports intention-focused digital activity. These insights informed the design goals of our system.
  • We present the design and implementation of the Intent Assistant (INA), an AI assistant that clarifies user intentions, detects context-dependent distractions and provides timely, non-intrusive interventions. We develop and publicly release the IntentionBench dataset, and use it to evaluate INA ’s distraction detection capability.
  • We validate INA in a three-week in-the-wild deployment with 22 participants, demonstrating its effectiveness in reducing off-task behavior, enhancing focus, and surfacing practical challenges for further improvement of intention-aligned assistance.
Together, these contributions advance the design of context-aware, ambient digital assistants by demonstrating how AI systems can pioneer adaptive, intention-aligned collaboration in everyday computing.
2. Related Work
Below we review three areas of related work: digital self-control tools, AI for digital well-being, and proactive assistants.
2.1. Digital Self-Control Tools
Digital self-control tools (DSCT) attempt to help people align their technology use with personal goals (Ko et al., 2015b; Roffarello and De Russis, 2023; Collins et al., 2014; Ahmed et al., 2014). Many prior studies have approached building DSCT with rule-based programs, such as allowing individuals or families to monitor usage against self-set limits (Hiniker et al., 2016; Ko et al., 2015a) or locking a device once a limit is exceeded (Kim et al., 2019). However, the rule-based DSCTs exhibit limitations, such as requiring manual configuration and a lack of adaptive interventions (Roffarello and De Russis, 2023). For example, a user might block YouTube with such a tool (because the website is typically a source of distraction) but then find the tool too inflexible to allow usage of YouTube for work tasks like learning about HCI (in contrast to the flexibility of INA illustrated in Figure 2). To address this, several studies attempt greater flexibility, such as varying the intensity or timing of interventions based on sensed mobile data like app usage history (Kovacs et al., 2018), allowing the relaxation of self-set rules like app blocks or time limits (Kovacs et al., 2021), and encouraging users to pause and reflect before engaging in potentially undesired actions (Cox et al., 2016). In this work, we propose a novel approach to building effective DSCTs: employing the remarkable capabilities of modern AI to detect the critical timing when a user’s activity diverges from their original intention and then providing timely, context-aware reminders.
2.2. AI for Digital Focus
Recently, various efforts have been made to leverage AI to reduce distraction. One of the attempts is building a prediction model of users’ off-task behavior while using digital devices from multi-modal data input, including app usage and screenshot (Chen et al., 2023). Another study reduced excessive smartphone use by locking the smartphone screen when overuse is detected. To unlock the screen, users are required to type specific digits, prompting a moment of reflection before continuing app usage. (Orzikulova et al., 2024) Other recent systems interact with users and improve their focus by utilizing large language models (LLMs) or conversational agents. For example, MindShift (Wu et al., 2024) intervenes when users attempt to open a blacklisted app by generating persuasive messages that discourage such usage. Also, StayFocused (Li et al., 2024) presents a specialized supportive application that blocks phone usage during a focus session, while providing reflective questions when the user tries to use it. While these approaches demonstrate the potential power of LLMs, their use of LLM is limited to message generation, and their detection of off-task behavior relies upon brittle rules. In contrast, our work aims to build a system that is context-aware in both its interactive messages and its off-task detection function, creating a deeply personalized experience.
2.3. Proactive Assistants
Proactive assistants are those that act without explicit user requests, shifting interactions from user-invoked to system-initiated while helping users follow through on their goals. Across many agent platforms, including robots (Darvish et al., 2025), mobile devices (Rawles et al., 2025; Lee et al., 2024b), and wearable devices (Fang et al., 2025), proactive assistants have been developed to provide various benefits to users, such as delivering timely, context-aware interventions and nudging users toward positive behavior change. For example, the Mirai system (Fang et al., 2025) provides contextually informed nudging, using a wearable camera and voice analysis to help users achieve pre-set goals such as “having a healthy diet” or “focusing on work.” However, the study primarily demonstrated the system’s technical feasibility through scenarios, rather than verifying its effectiveness on behavior change among real users with diverse goals. A recent study (Chen et al., 2025) introduced an LLM coding assistant that proactively provides suggestions to users when predefined, rule-based conditions are met. Building upon the tradition of these prior studies, we introduce a novel proactive assistant to promote intentional digital activities in general on-screen tasks.
3. Formative Study
We conducted formative interviews with individuals who use personal computers for long periods, with two goals: (1) to examine everyday digital activity and identify gaps between users’ intentions and their actual behavior, and (2) to explore the desirability of intelligent, context-aware interventions for supporting intentional digital activity. We describe the study procedure (Section 3.1), present key findings (Section 3.2), and outline the resulting design goals (Section 3.3).
3.1. Participants and Procedure
We recruited 11 participants (3 undergraduates, 4 graduate students, 4 office workers; ages 20–30, average age 25.6; 7 male, 4 female), whom we refer to as F1-F11. All participants reported using personal computers for at least four hours daily, providing a solid basis for examining discrepancies between intended and actual computer use. Eight participants had prior experience with productivity tools such as blocking or time-tracking applications, offering insights into the limitations of widely used conventional systems.
We conducted interviews focusing on three areas. First, we examined participants’ digital intentions: the goals they set for computer use, the strategies they employed to pursue these goals, and the patterns of deviation that emerged. Second, we investigated their experiences with existing productivity tools, including perceived benefits, limitations, and reasons for discontinuation. Third, we asked participants to envision an intelligent, context-aware assistant that could better support intentional digital activity, focusing on desired features, intervention styles, and timing. All interviews were audio-recorded with prior consent, manually transcribed, and analyzed using thematic coding. Participation was voluntary and uncompensated.
3.2. Findings
Ambiguous intentions and context-dependent distractions
Participants were asked about their typical focus periods and how they would set intentions when carrying out tasks on digital devices, providing up to three examples. Intentions were often expressed in high-level terms (e.g., “coding”, “reading papers”, “finish debugging”; F1, F3, F9) rather than in detailed statements. These broad expressions often encompassed multiple subtasks (e.g., searching for resources, implementing) and context-specific exceptions, making precise interpretation difficult.
Focus periods typically lasted 30–60 minutes but were often interrupted by distractions. Such distractions were commonly triggered by workflow stalls (e.g., reaching a dead end while debugging), boredom, or application notifications (e.g., email). Importantly, these patterns were highly context dependent: the same applications could serve as either task-related tools or sources of distraction depending on use. For example, YouTube enables both studying course videos and casual browsing, and email allows both on- and off-task communication.
Limitations of existing tools
Seven of the eight participants who had prior experience with productivity tools expressed dissatisfaction (F1–F4, F8, F10–F11), noting that these systems rely heavily on static, rule-based methods and therefore fail to account for context. As a result, users often had to manually disable blocking features to perform legitimate activities (e.g., accessing YouTube to study English, F11), which exposed them to additional distractions. Participant F11 described this frustration: “I had to forcibly disable the blocking program for studying purposes, only to get distracted and waste more time”. Such experiences underscore the limitations of static rules and point to the need for more intelligent, context-aware approaches.
Envisioning context-aware assistance
Participants described what an envisioned assistant for supporting intentional activity might look like. They preferred an assistant that could understand the user’s intention context, adapt flexibly to different situations, and intervene promptly when unintentional deviation from the stated intention was detected (F1, F9, F11). Rather than rigid blocking mechanisms, which were often perceived as intrusive and compulsory, participants expressed a strong preference for notification-based interventions that respected user autonomy. They also emphasized the importance of adjustable controls, such as fine-tuning how frequently the system should intervene or how persistent notifications should be (F3, F5, F6).
Tone and style of interventions were another recurring theme. While most participants preferred a gentle and polite tone for notifications (F4, F7), some considered friendly or more direct wording effective depending on the situation. Beyond distraction management, participants also envisioned interventions that could enhance productivity, such as offering praise or encouragement (F8), assisting with planning (F10), or providing intention-related suggestions when encountering obstacles (F8, F9). Taken together, these perspectives portray an assistant that is immediate in response, gentle in manner, and flexible in approach, not a strict enforcer but instead collaborative.
3.3. Design Goals
Based on the formative study, we derive four key design goals (DGs) to guide the development of a context-aware assistant.
  • DG1: Intention Understanding. The system must first enable users to articulate their intentions. Because such intentions are often abstract and open to multiple interpretations, the system should clarify them through interaction. Clearer intentions allow the system to anticipate on-task actions more effectively and minimize interventions that are based on incorrect detection of distraction.
  • DG2: Context-Aware Distraction Detection. The system should understand the user activity flexibly, going beyond simple blacklist-based rules. To this end, it must intelligently analyze contextual cues (such as on-screen content) in relation to user intention to understand their activity continuously and adaptively.
  • DG3: Timely and Gentle Interventions. When a distraction is detected, the system should intervene promptly to help guide the user back to their original intention. Interventions should be polite and dismissible notifications rather than enforced restrictions. The system may also reinforce positive behaviors by offering praise or encouragement when the user returns to their intended activity.
  • DG4: Feedback-Driven Refinement.
    Given that any contemporary system will detect distractionimperfectly, the user should be able to provide feedback on its detection to improve its future detection accuracy. This feedback can be used not only to improve general detection but also user-specific detection, for cases when the correct judgment of distraction in one context depends on the specific user.
Figure 2. Overview of INA. (1) The process begins with a user’s initial, often abstract, intention. (2) The system engages in an LLM Q&A interaction to (3) establish a clarified user intention. (4) An LLM detector then analyzes on-screen activity, using a distraction score to classify the user’s state as on-task or off-task. (5) Based on this state, the system delivers a gentle nudge or positive reinforcement. (6) The user can provide feedback on this intervention, which (7) the LLM uses to refine its model, improving detection accuracy for future interactions. :
4. Intent Assistant
In this section, we introduce the Intent Assistant (INA), an AI assistant that supports users by helping them accomplish their intended screen-based tasks. Our key idea is to build a context-aware system around a large language model (LLM), drawing on the LLM’s common-sense reasoning capabilities. As shown in Figure 2, INA consists of four main components, each designed to address the design goals (DGs) presented in Section 3.3. Each session begins when the user enters their intention as text and ends when the user presses the stop button. From this input, INA asks a few clarifying questions regarding their intention (DG1). The system continuously monitors the user’s on-screen activity to detect when they diverge from the stated intention (DG2). After detecting such distractions, INA intervenes with gentle, dismissible notifications (DG3). Over time, the system incorporates user feedback on the correctness of its detections to improve its detection accuracy (DG4). In the following subsections, we describe each component in detail.
4.1. Eliciting User Intentions
A session with INA begins when the user enters their intention as text. However, in our pilot study, we observed that user inputs vary widely, ranging from abstract statements (e.g., “study”) to very specific ones (e.g., “study HCI concepts in YouTube lectures”). While specific inputs make it easier for the system to relate user behaviors to their intentions, abstract inputs make this process more difficult.
To address the abstract nature of user intentions, we incorporate a clarification process with short Q&A interaction using the LLM. Given an initial intention, the LLM generates a clarifying question to better understand the user’s goal. If the user responds, the LLM may generate a follow-up question to further refine the expressed intention. For example, when the initial input is “study”, the LLM may first ask “What subject are you planning to study, such as math, history, or a language?”, narrowing the scope to a specific domain. If the user answers, “I will study HCI concepts”, the LLM can then follow up with “What tools or resources will you use, such as textbooks or online courses?” shifting the focus from what to how the user intends to study. The LLM asks two questions and allows the user to stop the dialogue at any time. Please refer to the Appendix A.1 for the details of this procedure.
4.2. Monitoring User Behavior to Detect Distraction
Given a user’s clarified intention from the Q&A process, our system continuously monitors on-screen activities to evaluate whether users remain focused on their stated goals. A central challenge in this process is that distractions are not always identifiable from surface-level signals such as URLs, application names, or keywords. For example, two users may both be watching videos on YouTube: for one, the activity directly supports their goal of studying HCI concepts, while for the other, it constitutes an unrelated diversion.
To address this challenge, we leverage LLMs to assess the semantic alignment between user activities and their stated intentions. Specifically, the LLM is prompted to evaluate alignment and generate a distraction score—a score in $[0,1]$ that ranges from perfect alignment (0) to complete misalignment (1)—based on three input signals:
  • Intention: The clarified intention that was generated by the LLM from both the original intention and the Q&A process
  • Current screenshot: A single screenshot of the user’s current screen
  • Application metadata: Information about the application in focus, including its title and, if the application is a browser, the current URL [^1] This application metadata is especially important since certain contents (e.g., streaming videos) may not appear in screenshots because the platform prevents capturing them.
By integrating these inputs, the LLM assesses how observed behavior cues (e.g., text, images, UI elements, application functions) relate to the user’s intention, rather than relying solely on keywords or application categories. The output format encourages the model to generate a brief chain-of-thought rationale before producing a distraction score, which improves both accuracy and interpretability (see Appendix A.2 for the full prompt). The resulting distraction score provides the quantitative basis for the subsequent process, i.e., judging whether the user is distracted and delivering notifications for gentle guidance.
4.3. Supporting Users According to their Distraction State
Given a distraction score, INA classifies the distraction state of the user as on-task if the distraction score is less than 0.5—indicating that the user is currently aligned with their stated intention—or as off-task otherwise, meaning they are distracted. Upon transition from one distraction state to another, a notification is issued if and only if a change between these states is sustained for a certain duration (4 seconds), rather than responding to brief or accidental shifts. Specifically upon an on-$\rightarrowoff-task$ transition, INA delivers a gentle nudge that invites the user to return to their stated intention. On the other hand, upon an off-$\rightarrowon-task$ transition, the system provides praise to reinforce the user’s regained focus. If the user remains in the off-task state, the nudge is repeated at fixed intervals (every 30 seconds), a compromise chosen to remain noticeable without being disruptive, whereas no further notifications are issued while the user continues to stay on-task.
Notifications take the specific form of pop-up messages that automatically disappear after a short duration (see Section 4.5 for illustration). Furthermore, the off-task messages are generated by the LLM in INA to be polite questions rather than commands (see Appendix A.3 for the full prompts). For example, when distraction is detected, the system may ask, “It seems your attention is on ‘online shopping’. Shall we restart with ‘studying HCI’?”. A return to focus is acknowledged with short praise such as “You are focused on ‘watching lectures on YouTube’. Great work!”.
Figure 3. (1-2) The user inputs an intention and presses ‘Set’ to initiate a session. (3-5) A chat interface is used for a brief Q&A clarification process, where the user responds to the LLM’s questions. (6) The user can also press ‘Start’ to skip this clarification and begin the session immediately. (7) During the session, INA displays the user’s current state with colored notifications: green for on-task and red for off-task. (8-9) The user can hover over a notification to provide feedback, marking it as correct or incorrect and optionally adding a reason. (10) The ‘Stop’ button can be pressed at any time to end the session. :
4.4. Improving Accuracy Through User Interaction
To improve detection accuracy, INA incorporates a feedback process, an interaction loop that refines distraction detection based on each user’s digital habits and situational context. After each notification, users can provide quick feedback by marking the system’s distraction detection as correct or incorrect (see Section 4.5 for the depiction of the implemented UI). This interface is designed to enable users to provide feedback quickly and easily, with a single click.
Once feedback is collected, INA leverages its underlying LLM to analyze misclassified cases and refine its future evaluation of the semantic alignment between user activities and stated intentions, as described in Section 4.3. Specifically, when a notification is marked as incorrect, the system records the surrounding context (e.g., screenshot, active application or URL, and the distraction score) and prompts the LLM to analyze why the detection was inappropriate. The LLM then generates a short refinement that is appended to the prompt for the remainder of the session, enabling more accurate alignment measurement in similar future situations.
For example, if a user marks as incorrect a notification raised while they were searching for learning resources on YouTube, the LLM is prompted to analyze the activity to find relevance that the browsing activity was indeed “to find HCI relevant videos”. Accordingly, a short refinement, such as “output lower score when detecting activity - YouTube search results page for ‘HCI lecture’,” is included in the prompt as additional input for the scoring process. Through repeated interactions, these corrections accumulate and progressively reduce false alerts. We evaluate this feature empirically in Section 5 and provide the full prompt used for refinement in Appendix A.4.
4.5. System Implementation
We illustrate the user interface of INA in Figure 3. Users first type their intention into a text field and press the Set button to initiate a session. A brief clarification process then takes place in a small chat interface, where blue messages represent questions generated by the LLM and white messages represent user responses. If users consider their input already specific enough, they can skip this step and start the session immediately by pressing the Start button. Every two seconds during the session, INA classifies whether the user’s activity is on-task or off-task. Following the notification rule explained in Section 4.3, INA might deliver a gentle notification to the user. The message in the notification is also visible through a box in the app window (green when the user is on-task or red when the user is off-task). Finally, users can provide feedback on notifications by hovering over them, marking whether they are correct or incorrect, and optionally adding a short explanation. Each feedback result is appended to subsequent input prompts and retained for a day, preventing unnecessarily long prompts and thereby reducing the cost.
The implementation of INA was developed in Python. The client is a native macOS application built with PyQt to provide a standardized experimental environment, and the backend is implemented with FastAPI and deployed on Google Cloud Platform. LLM queries are processed through Gemini 2.0-Flash with a temperature of 0.1, while fixing other parameters to the default. To protect sensitive data, screenshots are masked using Cloud DLP (Data Loss Prevention) before upload, with images stored in Cloud Storage and corresponding metadata and event logs maintained in Firestore. Further technical details about implementation of INA are provided in Appendix A.5, and privacy and security considerations are discussed in Appendix E.
5. Evaluation of Distraction Detection
In this section, we evaluate the distraction detection capability of INA using a dataset constructed to mimic user workflows, named IntentionBench, enabling rigorous validation under static yet realistic conditions.
5.1. IntentionBench
Figure 4. An illustration of the IntentionBench construction process. We first collect focused sessions, where each session corresponds to executing a distinct task instruction (e.g., “Study HCI”). Then, two focused sessions are combined to synthesize a mixed session by concatenating and reordering their segments. The user intention is inherited from one session (on-task), while the other constitutes off-task segments, mimicking natural transition points of user workflows.:
Ideally, real-world usage data from end users would provide the most direct validation of system performance. However, such data present two major challenges: (1) collecting them at scale is costly and time-consuming, and (2) transitions from on-task to off-task -which are the points at which INA’s most critical notifications are triggered—would comprise a small minority of the user data. This imbalance makes it difficult to rigorously evaluate detection accuracy at these critical points.
To address these limitations, we constructed IntentionBench, a novel dataset based on real activities performed by the authors. We first collected 50 focused sessions spanning diverse scenarios (14 applications and 32 websites). Each focused session was generated by executing a distinct task instruction (e.g., “Plan a winter trip abroad”) with two collectors acting as users and capturing screenshots every second. Before executing each instruction, they also performed a Q&A interaction to clarify their plan for the activity execution (e.g., trip location or websites to be used) to simulate the intention clarification process (Section 4.1). Then, each collected focused session was divided into smaller segments, with boundaries defined by natural transitions such as switching applications or navigating to new websites.
From these focused sessions, mixed sessions were then synthesized, as shown in Figure 4. Specifically, mixed sessions were created from randomly sampling two focused sessions, concatenating them, and then randomly reordering their segments to generate activity transitions. The instruction from the first focused session serves as the user intention: segments from this focused session were labeled on-task, while segments from the second were labeled off-task. This mixing process balanced the proportion of intended and unintended activity, added numerous transitions to off-task status, and also retained the transition points that occurred during the real human-executed behavior. As a result, IntentionBench comprises 350 mixed sessions, each approximately 13 minutes on average in duration. IntentionBench totals 138,803 data points (each consisting of a screen capture, the user intention, and, when available, results of clarification corresponding to the user intention), covering approximately 77 hours of activity. Further details are provided in Appendix B. We will release IntentionBench publicly.
Clarification Feedback Accuracy Precision Recall F1
0.805 0.897 0.629 0.739
0.871 0.949 0.748 0.836
0.845 0.959 0.677 0.794
0.878 0.959 0.755 0.845
Table 1. Effectiveness of INA on IntentionBench. The check mark (✓) indicates whether clarification or feedback is included; the best performance is in bold. Employing both features shows the best balance. The bottom row is our deployed configuration.
5.2. Evaluation Results and Analysis
We evaluate the performance of INA in detecting distraction using IntentionBench. For each data point, INA is prompted with the user’s stated intention and the corresponding screenshot, and produces a distraction score between $0.0$ and $1.0$. Distraction scores below $0.5$ are classified as on-task, and scores of $0.5$ or above are considered off-task. The predicted labels are compared against the on-task/off-task labels generated along with the mixed sessions, and performance is assessed using four standard metrics to capture complementary aspects of performance: accuracy, precision, recall, and F1-score.
Table 1 summarizes detection performance with IntentionBench for our full system (bottom row) and under ablations: removing one or both of clarification and feedback features to analyze their contributions. [^2] For feedback conditions, user feedback is simulated with a simple heuristic that corrects every false positive, forming an upper bound. Then, evaluation proceeds sequentially through each mixed session to ensure that corrections are incorporated in temporal order. See Appendix B.2 for more details). Without either of these two interactive features, the system achieves an accuracy of 0.805 and an F1-score of 0.739. Adding the results from clarification interaction yields substantial gains (accuracy = 0.871, F1 = 0.836). Incorporating feedback improves accuracy and F1 (0.845 and 0.794, respectively) over INA with neither feature. Finally, combining both clarification and feedback produces the best overall performance (accuracy = 0.878, F1 = 0.845). Therefore, whereas Q&A creates larger improvements than feedback—except with respect to precision—the best version of INA requires both. We further validate INA on a dataset curated from real-world deployment records (Appendix B.3), where it attains an accuracy of 0.899. This additional result demonstrates that INA remains effective under the noisy and ambiguous conditions of real-world usage.
Figure 5. Overall procedure of our three-week, within-subjects field study. Participants attended a 30-minute online orientation session before the study. Participants used one application type on their computers each week. The assignment of each application type—INA, logging only, or simple reminder—was based on a random order. At the end of each week, participants completed a post-survey and a 10-20 minute semi-structured video interview.:
6. User Study Design
We conduct a three-week long, within-subjects field study to evaluate whether INA effectively supports intentional digital activity. Our study investigates the following research questions:
  • RQ1: Do users carry out their intended tasks with greater focus when using INA?
  • RQ2: What is the overall user experience with INA?
6.1. Participants
We recruited participants through social media platforms and online communities of university students and job-seekers. In total, 81 individuals filled out a pre-survey that collected demographic information and assessed computer usage habits (see Appendix C.1 for details). [^3] The pre-survey also contained a Short Self-Regulation Questionnaire (SSRQ; (Carey et al., 2004; Brown et al., 1999)), but the answers were not used in recruiting nor analyses because they were later deemed unrelated to our research questions. Among these applicants, we selected those who meet these three criteria: (a) uses a MacBook (for application compatibility), (b) are not employed at a corporation (to reduce data security concerns; eligible participants are university students or job seekers), and (c) report at least a moderate level of digital distraction, ensuring that the study focuses on a population with needs for the intervention.
Our study began with 24 participants, of whom 2 voluntarily withdrew during deployment. Consequently, we analyze data from 22 participants (14 women, 8 men), aged 19–39: 9 aged 19–24 (40.9%), 8 aged 25–29 (36.4%), and 5 aged 30–39 (22.7%). The average self-reported computer use time is 5.6 (SD=3.02) hours per day. Participants were compensated at approximately 1 USD per hour of program use, up to a maximum of 72 USD, provided via a gift card. The study protocol was approved by an Institutional Review Board.
6.2. Baseline Systems: Logging only and Simple Reminder Applications
To evaluate the efficacy of INA, we develop two baseline applications: logging only and simple reminder applications. The logging only application records desktop activity without requesting users’ intentions or sending notifications. The simple reminder application asks users to state their intention at the start of each session, reminds them of this intention every 25 minutes, and displays the current user-stated intention in the status bar. This application was designed to isolate the effect of simply articulating and periodically reviewing one’s intention (see Appendix C.2 for more details). In contrast, INA collects the user’s intention at the beginning of each session with clarifying questions, monitors user activity for off-task behavior, and delivers context-aware notifications when users become distracted. In both the simple reminder and INA, users rated how well their behavior matched their stated intention on a five-point scale at the end of each session. Across all three applications, screenshots, names of active applications, and the URL of the window that is in focus (i.e., front-most on the screen) were logged every two seconds after participants pressed the start button. All logs were anonymized to ensure participant privacy (see Appendix E for details on our data protection methods).
6.3. Procedure
Figure 5 visualizes the overall procedure of the study. The study was conducted over three weeks as a within-subjects field deployment. Initially, participants attended a 30-minute online orientation session prior to deployment. In the orientation, we explained the overall goal of the study (i.e., to evaluate an AI agent that supports focused and intentional digital device use), data-collection policies, and instructions for using the three applications. While we briefly mentioned the differences among the applications, we avoided detailed descriptions to reduce differences in participants’ expectations of each application. [^4] We discuss the potential limitation associated with this orientation in Section 8.2.
During deployment, participants used all three applications described in Section 6.2, each for seven days using their computers. The order of applications was randomized for each participant to address ordering effects. Application names were masked and presented only as color labels (e.g., Purple, Blue, Orange) to reduce potential bias. Participants were instructed to use their computers for at least two hours per day with the assigned application running to ensure sufficient data collection. At the end of each week, participants completed a weekly post-survey. On the following day, they participated in a 10-20 minute semi-structured video interview.
6.4. Measures
To evaluate whether participants carried out their intended tasks with greater focus with INA (RQ1), we used the following measures:
  • LLM-estimated off-task ratio: The LLM-estimated off-task ratio is defined as the proportion of data points (sampled every 2 seconds) classified as off-task based on the distraction score evaluated by an LLM, as introduced in Section 4.2. This metric is available only in simple reminder and INA, as it requires knowledge of user intentions. A lower ratio indicates that participants spent more time aligned with their stated intentions.
  • Intention alignment rating: The intention alignment rating is the self-reported score collected at the end of each session (see Appendix C.2 for the rating interface), where participants answered the following question: “How well did your activity align with your intention?” on a 5-point Likert scale (1 = not aligned, 5 = very well aligned). This metric is also available only in simple reminder and INA, as it requires stated user intentions.
  • Focused immersion scale: The focused immersion scale measures the level of participants’ concentration during their digital activities. We modified the Focused Immersion Scale (Agarwal and Karahanna, 2000) to ask about computer experience (i.e., “While using the Web” to “While using the computer”), using a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree). Participants answered this survey as part of the post-survey that was completed at the end of each one-week deployment period. See Appendix C.3 for more details.
To assess users’ overall experiences with the three applications (RQ2), the weekly post-survey also included the following questions:
  • User experience questions: We adapted and expanded upon the questionnaire from GPTCoach (Jörke et al., 2025). Specifically, we modified some of the original questions to better fit our research context and developed several new items to assess aspects unique to our system, such as message effectiveness and workflow disruption. A factor analysis of the initial 15 questions resulted in 3 final scales, constructed from 13 of the 15: Support, a 7-item scale measuring how participants perceived that the application supports the user’s adherence to their intentions (Cronbach’s $\alpha$ = .81); Message Effectiveness, a 5-item scale assessing the perceived effectiveness of system notifications (Cronbach’s $\alpha$ = .78); and Workflow disruption, a single-item scale gauging negative impacts such as workflow interruptions. All questions were on a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree). The complete list of all survey questions, along with the rationale for excluding two questions, is available in the Appendix C.3.
We also conducted weekly semi-structured interviews to gain in-depth, qualitative insights about participants’ experiences using the applications (RQ1 & RQ2). In each of these 20-minute sessions, participants were asked to describe their overall experience in the past week, provide anecdotes from their usage of the application for the week, and share their perceptions of the application’s impact on them. The complete set of user experience surveys and interview questions is presented in Appendix C.3.
6.5. Methods of Analysis
Statistical Analysis
All participants experienced the three different applications across multiple sessions. To account for repeated measures within participants, we analyzed the experimental data using a linear mixed-effects model (Lindstrom and Bates, 1988). We model each measure by treating the application type as a fixed effect and the participant as a random intercept. Formally, this is specified as $\mathrm{measure}\sim\mathrm{program}$ + $\left(1\mid\mathrm{user}\right)$. When comparing the three applications, pairwise contrasts are obtained from estimated marginal means, and the $p$-values are adjusted using the Benjamini–Hochberg false discovery rate procedure (Benjamini and Hochberg, 1995). For measures involving only two applications (simple reminder vs. INA), we directly report the $p$-value from the linear mixed model without adjustment.
Interview Analysis
Interviews were transcribed and coded through a thematic analysis. Two researchers collaboratively constructed and iteratively refined the codebook via inductive open coding, following three principles: (1) codes were assigned based on the semantic surface of utterances to minimize interpretive bias, (2) multiple codes could be applied to a single utterance, and (3) themes were counted at most once per participant (i.e., respondent-level binary aggregation). Appendix D.4 provides the detailed coding procedures and intercoder agreement statistics, while Table 2 presents the full codebook along with summarized results.
(a) LLM-estimated off-task ratio ($\downarrow$)
(b) Intention alignment rating ($\uparrow$)
Figure 6. (a) Mean LLM-estimated off-task ratio. (b) Mean intention alignment rating based on end-of-session self-reports. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are more favorable.:
7. User Study Findings
In this section, we present findings from our three-week user study. In Section 7.1, we describe overall system usage and user feedback. Section 7.2 reports the impact of INA on participants’ intentional activity. We then summarize results from the user experience survey in Section 7.3, followed by qualitative insights from post-study interviews in Section 7.4.
7.1. System Usage
Over the three-week study period, each participant used applications on average for 2.9 hours per day, amounting to roughly 20.6 hours per application across the three one-week conditions. This resulted in a total of 1,786 sessions, amounting to 1,360 hours of usage data and 2,449,690 system logs (screen images, app log information). On average, each session lasted 45.7 minutes, and sessions shorter than one minute were excluded from the analysis. Across simple reminder and INA sessions, participants entered 1,248 distinct intentions and provided 1,071 intention alignment ratings (85.8%) [^5] Survey data was not collected if the session ended prematurely due to application bugs or computer shutdowns. at the end of their sessions.
Furthermore, we examined participants’ feedback to INA. Out of 1,786 sessions, 19 participants provided explicit feedback during 90 sessions, resulting in 161 feedback instances. The majority of feedback was positive (111 ‘correct’, 68.9%), while 50 instances were negative (‘incorrect’, 31.1%). Notably, 42 feedback events contained free-form messages from users, allowing us to gain a deeper understanding of their experiences. Typical positive feedback messages expressed recognition and surprise (e.g., “accurate” and “you caught me”), while negative messages often explained the relevance of activities to their intentions or guided how the system should have worked.
7.2. Impact of INA on Intended Activity and User Focus
As shown in Figure 6, for the two applications that collect user intentions, INA not only reduces off-task time during task execution but also increases users’ reported alignment between their digital activity and their stated intention. The LLM-estimated off-task ratio [^6] Our technical evaluation in Section 5 provides evidence of its accuracy. is significantly lower with INA than with simple reminder (0.104 vs. 0.166, $p<.001$). Users’ self-reports reinforce this finding: intention alignment ratings—users’ perceived alignment between their activities and stated intentions—are higher under INA (4.44 vs. 4.23, $p<.001$). These results together indicate that users exhibited increased intentional activity with INA.
To more directly analyze the relationship between intention alignment ratings and LLM-estimated off-task ratios at the session level (see Appendix D.2), we calculate their Pearson correlation and find it to be weakly yet significantly negative (r=0.196,p<0.01r=-0.196,p<0.01), indicating that lower off-task time is associated with higher alignment ratings, as expected. Further analysis across activity types is reported in Appendix D.1.
Users also self-assessed their concentration over the previous week via the focused immersion scale in the post-survey. As shown in Figure 7(a), INA achieved higher scores than the baselines (3.74 vs. 3.34 for simple reminder, $p=.0449$; and 3.74 vs. 2.90 for logging only, $p=.0003$), providing further evidence that INA aided users’ focus. The focused immersion score further revealed significant differences across all three applications: even simple reminder outperformed logging only (3.34 vs. 2.90, $p=.0440$), suggesting that simply entering and being reminded of one’s intention already has a positive impact on focus, while INA, with its context-aware interventions, produced the largest improvement. In the interviews, all participants except one reported positive experiences with INA, noting improvements in work efficiency, such as “It [INA] helped me focus more quickly, so my work efficiency improved” (P15).
7.3. User Experience Survey
(a) Focused Immersion Score ($\uparrow$)
(b) User experience survey scores ($\uparrow$)
Figure 7. (a) Mean focused immersion scores, a self-reported measure of concentration taken during the weekly post-survey. (b) Mean user experience survey scores, computed by averaging responses within categories: Support , Message Effectiveness, Workflow Disruption. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. ($\uparrow$) indicates that higher values are more favorable. Additional details are in Appendix D.3, including the scores for each question.:
Theme Code Definition logging only simple reminder INA
Pos Neu Neg Pos Neu Neg Pos Neu Neg
Effect C1. Achievement Task completion (possibly, partial) 13.6 18.2 63.6 72.7 18.2 4.6 86.4 4.6 0.0
C2. Distraction reduction Reduction in off-task & Success in return 0.0 4.6 22.7 31.8 4.6 13.6 63.6 4.6 0.0
C3. Focus Sense of focus/accomplishment 13.6 18.2 9.1 36.4 0.0 0.0 31.8 0.0 0.0
C4. User reaction Acceptance/ignoring patterns to notifications 31.8 22.7 18.2 63.6 18.2 13.6
Notifications C5. Context Contextual fit or misclassification 18.2 31.8 22.7 59.1 9.1 22.7
C6. Timing Timing and frequency of intervention 36.4 13.6 22.7 31.8 18.2 31.8
C7. Impression Tone/length; intrusiveness 36.4 22.7 13.6 59.1 9.1 9.1
Support C8. Care Feeling of being cared/managed 18.2 0.0 13.6 22.7 9.1 9.1 50.0 0.0 0.0
C9. Motivation Feeling of being motivated 0.0 9.1 27.3 13.6 31.8 9.1 31.7 22.7 4.5
C10. Closeness Companionship/familiarity/feeling as a peer 0.0 0.0 0.0 0.0 0.0 4.6 22.7 0.0 0.0
C11. Reflection Opportunities for reflection/introspection 13.6 9.1 4.6 27.3 22.7 0.0 40.9 4.6 13.6
Inputting intention C12. Effect of inputting Inputting itself enhanced intentional activity 77.3 0.0 0.0 18.2 4.6 0.0
C13. Convenience (input) Inputting process felt effortless/convenient 36.4 13.6 31.8 9.1 9.1 13.6
Interactive features C14. Effect of clarification Clarification supported intentional activity 40.9 9.1 4.6
C15. Convenience (clarification) Burden of detailed dialogues/inputs 27.3 13.6 36.4
C16. Effect of feedback Improvement in notification due to feedback 27.3 36.4 27.3
Adaptation C17. Offline spillover Spillover to offline behavior/habits 9.1 4.6 13.6 36.4 18.2 9.1 45.5 27.3 9.1
C18. Habit/strategy change Behavioral changes (e.g., break-blocks, strategies) 9.1 4.6 18.2 9.1 0.0 9.1 13.6 13.6 0.0
C19. Long-term trends Early vs. later stage adaptation, residual effects 18.2 27.3 13.6 36.4 13.6 27.3 31.8 22.7 13.6
Concerns & Hindering C20. Privacy concerns Repeated concerns over personal data 40.9 9.1 31.8 59.1 22.7 9.1 54.6 9.1 22.7
C21. Workflow disruption/load Interventions breaking flow or adding load 40.9 4.6 0.0 27.3 13.6 9.1 4.6 13.6 40.9
C22. Distrust Perceived distrust or ineffectiveness 0.0 0.0 68.2 4.6 4.6 4.6 0.0 0.0 0.0
Willingness to adopt C23. Reuse intention Willingness to use if released 9.1 4.6 77.3 50.0 13.6 22.7 63.6 13.6 13.6
C24. Willingness to pay Positive willingness to pay 0.0 0.0 9.1 13.6 4.6 22.7 27.3 31.8 4.6
Table 2. Prevalence of interview codes by theme and applications (simple reminder, logging only, and INA). All values are expressed as percentages. For each code, the numerator counts the number of participants who mentioned the code at least once (with a maximum of one count per participant), and the denominator is the total number of participants, $N=22$. Cells denoted by “–” indicate that the information is not applicable. Each code was further classified as positive (Pos), neutral (Neu), and negative (Neg).
Figure 7(b) summarizes users’ overall experiences across three categories from the weekly post-survey. In the Support category, participants rated INA as significantly more effective than logging only in supporting intentional digital activities (3.75 vs. 2.75, $p<.001$). Compared to simple reminder, INA also received higher ratings for supporting intentional digital activities, although this difference did not reach statistical significance (3.75 vs. 3.56, $p=.267$). As shown in the Message Effectiveness category, users identified the context-aware, timely notifications of INA as a key factor underlying its benefits, as evidenced by higher scores for INA than for simple reminder (3.74 vs. 3.24, $p=.013$). In contrast, INA received the lowest score in the Workflow Disruption category (lower scores indicate more disruption) compared to simple reminder (3.09 vs. 3.91, $p=.005$) and logging only (3.09 vs. 3.91, $p=.005$). This aligns with the main concern raised in interviews, where participants noted that INA could be bothersome due to excessive interruptions.
7.4. Interview Findings
We provide a detailed analysis of user experiences of 22 participants (P1–P22) with the three applications, as revealed during the interviews. See Table 2 for the interview codes (C1–C24) and the percentage of responses for each code. Text in square brackets is used to provide context and aid understanding of the quotation.
Timely, context-aware notifications of INA were recognized to be effective
Many participants pointed out that INA helped them achieve what they intended, reducing time spent on distractions (see C1 and C2 in Table 2). P14 portrayed this experience to a “car’s lane-keeping assist feature”, stating, “The moment I got distracted by YouTube while studying, a notification helped me return to my intended task”. The participants often noted that the core effectiveness of INA stemmed from its ability to deliver a timely, context-aware notification when they became distracted (see C5 in Table 2). They reported that alerts enabled them to immediately recognize they were off-task and quickly return to their original work. P3 commented on the impact of INA’s contextual understanding upon their activity, stating [INA understood] the research-relatedness of the dialogue contents when I was in Messenger”. Several participants also emphasized that the timely delivery of notifications was essential (see C6 in Table 2), as in “It [INA] notified me right away when I got sidetracked with email or YouTube during a meeting” (P4) and “the immediate notification allowed for a swift return to the original task” (P20).
In contrast, users of the simple reminder were only about half as likely to report a reduction in distraction as those using INA (see C2 in Table 2). Participants explained that this was because the simple reminder functioned more like a passive “post-it” (P10) on their screen rather than an active, context-aware notification. This lack of context-awareness was noted as a critical limitation of simple reminder. Many noted that because messages were static and did not adapt to their activity, notifications were ineffective and easily ignored. For instance, participants commented “Even if I was doing something different from my intention, no specific notification appeared, so I didn’t really pay attention” (P12), and “The messages didn’t really stand out, so I just overlooked them” (P13).
INA was perceived as a supportive and motivating assistant
We observed that some users felt a close connection with INA (see C8–C11 in Table 2). Indeed, P1 portrayed INA as secretary or a parent, suggesting a supportive and caretaking role. P2 also described INA as “somewhat interactive, almost like a mate”. P4 explained that INA “did not feel mechanical but rather like a one-on-one manager offering personalized support”. The effect of praise, especially when users were on-task, was particularly notable. For example, P1 noted that “When I received positive messages, I felt proud and recognized… compliments boosted my self-esteem”, and P7 stated “I strongly felt as if I was being cheered on”. On the other hand, the participants mentioned less on support of simple reminder and logging only compared to INA. P19 noted that “the messages felt like conveying obligations without a sense of warmth or companionship”.
The process of specifying the intention itself was reported to foster more deliberate and mindful digital behaviors
The positive influence of typing intentions was consistently observed (see C13 in Table 2). In both the INA and simple reminder groups, the act of physically inputting an intention helped participants gather their thoughts and plan their actions. As P1 and P5 noted, “Writing down my intention made me use the computer more deliberately, so I rarely got sidetracked”. P11 also added, “Inputting the goal feels like definitively deciding what I am going to do right now before proceeding”. These findings show that the essence of this process lies in the cognitive support of self-declaration, enabling users to think concretely about their intentions and clearly recognize the actions they need to take.
Moreover, participants explained that process of clarification of INA helped them specify vague ideas into more concrete plans (see C14 in Table 2). For example, P2 and P9 commented that the clarification process made them “clarify on my intentions [before carrying out tasks]. Specifically, P19 recalled, “When I entered my intention only as something like draft the IRB, it [INA] asked, ‘Which part will you be writing?’ As I answered it to write ‘the research purpose’ part, my initially rough intention became increasingly concrete, which I appreciated”.
INA also promoted benefits during users’ offline activities.
Several participants noted offline spillover from use of INA, mentioning the impact on their daily routines (see C17 in Table 2). P12 stated, “I often stayed up late using my computer at night, but since it curtailed my off-task time, I felt I could go to bed earlier”, mentioning positive effects of INA in their routines. P19 mentioned, “the frequency of me getting sidetracked has decreased even in my daily life”. We believe that designing long-term strategies to sustain this effect could be valuable future work. Similar benefits were also mentioned while using simple reminder, but with a smaller frequency than INA, while such effects were mentioned the least with logging only.
Downsides of INA, despite effectiveness
From the interviews, we also identified several concerns about INA, primarily related to workflow disruption, inconvenience, and data privacy. Some participants experienced workflow disruption from the system’s notifications (see C21 in Table 2). We acknowledge that notifications can inevitably feel bothersome since they interrupt what users are doing despite a supportive purpose, as in “I felt the alerts came too frequently and were somewhat distracting” (P12). To address such comments, we believe that improving distraction detection capability or allowing users to control the frequency by adjusting the distraction detection threshold can be a future direction.
The interactive components in INA also introduced a certain degree of burden for users, mainly because interactions required additional effort (see C15 & C16 in Table 2). For example, P12 added “I repeat [the Q&As for every start of a session] … it felt frustrating. I wondered, Why am I inputting this again?”. Also, the lack of generalization ability of LLMs related to feedback and refinement required the users to repeat the corrections, frustrating several participants (P5, P11).
Importantly, about half of the INA users reported concerns regarding the privacy of sensitive data (see C20 in Table 2), despite the presence of several safeguards such as anonymization before storing data. Some participants were aware of these safeguards, as in “I felt reassured about the privacy protection system” (P3). However, other participants expressed worries, particularly when the tasks involved sensitive information. For example, P17 stated “When dealing with payment-related work, I turned off the program”, and P13 mentioned “I felt cautious when signing up for memberships”. P9 expressed unease about the system recognizing messenger conversations, noting that “When chatting with others, it [INA] told me not to get distracted … which felt impressive but also a bit concerning”. These findings highlight the importance of not only establishing robust privacy safeguards but also communicating them transparently to users, especially in contexts involving highly sensitive information. Further, users might feel more at ease outside of an experimental setting like ours—in which researchers were explicitly collecting records of their activity—especially with a future version of INA that works with an on-device LLM, keeping their data only on their device.
Figure 8. An illustration of INA’s safety challenges when faced with harmful intentions. (1) A user states a harmful or unsafe intention, such as ”Hacking into a hospital server.” (2) As the user performs actions aligned with this goal (e.g., visiting a medical records page of a hospital portal, watching a tutorial on phishing emails), INA provides positive reinforcement, encouraging the dangerous activity. (3) When the user’s behavior deviates to a safe but unrelated topic (e.g., watching an ethics lesson), the system perceives this as a distraction and delivers a notification, prompting the user to return to the original harmful intention. Such behavior occurs even though the underlying LLM is expected to refuse harmful requests, and shows the need to introduce additional safety measures.:
8. Discussion
In this section, we summarize our findings, discuss limitations and future directions, and consider potential risks, particularly in supporting harmful user intentions.
8.1. Summary of Findings
We summarize our main findings from Section 7 in relation to our two research questions:
  • RQ1: Do users carry out their intended tasks with greater focus when using INA? INA effectively increased the proportion of participants’ activity that was spent on their intended tasks, with participants demonstrating reduced off-task time proportion and a higher match between their activity and intention (Section 7.2).
  • RQ2: What is the overall user experience with INA? In comparison to the other two applications, users perceived INA as more valuable for guiding their digital activities toward their intentions and experienced increased focus, but they also raised concerns about data privacy and excessive intervention (Sections 7.3 and 7.4).
Overall, these findings suggest that INA effectively helps users maintain focus on their intended tasks and is perceived as a valuable assistant in guiding their digital activities. At the same time, its success depends on addressing user concerns regarding privacy and the intrusiveness of interventions. This tension between interventions helping distraction yet also potentially causing distraction highlights important design considerations for future intention-support systems.
8.2. Limitations and Future Directions
On user experiences
Reflected in positive interview and higher user experience survey ratings for support and effectiveness, participants accepted INA as a focus-restoring aid but also reported fatigue due to taxing aspects of its user experience such as Q&As being required for every session, feedback records lasting only for a day, and workflow disruption from frequent notifications. To prevent such fatigue, future work could adopt adaptive policies that personalize intervention frequency, timing, and intensity, with either user-facing controls for baseline preferences (e.g., message tone, timing) or automatic preference learning from interaction data. We believe that such personalization would substantially improve user acceptance and effectiveness of the system.
Furthermore, while we extensively examined the effects of various system features, each system was tested with a single UI design. However, UI design itself can be a critical factor shaping user experience and engagement. An interesting direction of future work would be exploring alternative design variants to examine how UI differences influence user outcomes and to identify design principles that best support intention management.
Impact of being recorded
We acknowledge that participants’ awareness of being recorded may have influenced our study to some degree. Some participants reported that simply being recorded during the experiment gave them a sense of being monitored and managed, which in turn affected their behavior (P3, P17). While we attempted to control for this effect by including the logging only application, allowing us to separate the influence of mere-recording, future work could adopt alternative experimental designs that reduce users’ immediate sense of being part of an experiment.
Although we made strong efforts to anonymize the data and ensure careful handling, about half of the participants still expressed concerns regarding the treatment of sensitive information. These concerns suggest that technical protections alone may not be sufficient, and future work should explore building stronger privacy safeguards. For example, an on-device system that processes sensitive data locally, without transmitting it to external servers, could reduce the privacy risk.
Potential bias in the user study
During the orientation, participants were informed that one of the three applications included an AI agent, without specifying which role it occupied. Other minor differences and similarities, such as intention prompting and screenshot collections, were also introduced. To minimize potential bias and prevent participants from inferring features based on application names, the applications were color-coded from orientation through the study period. Nevertheless, we acknowledge that knowing one application involved an AI agent—as well as —and from that knowledge, perhaps assuming that the researchers preferred it— may have led participants to hold different attitudes towards and expectations for that version.
Risks of encouraging unsafe intentions
Similar to the risks posed by AI agents (Andriushchenko et al., 2025; Lee et al., 2024a, 2025), INA also carries the risk of reinforcing unsafe intentions. In Figure 8, we illustrate an example case where INA provides positive reinforcement to a harmful intention “hacking into a hospital server”. Surprisingly, INA did not refuse to provide positive feedback to such harmful intentions, even though one might normally expect the underlying LLM to have the capability to reject them. We hypothesize that this behavior persists even when common refusal mechanisms are in place, since the LLM is tasked with predicting a user’s intentions rather than directly executing the harmful request. To mitigate such outcomes, we expect that integrating external guardrail models (e.g., WildGuard (Han et al., 2024)) can effectively detect harmful intentions and prevent the system from providing encouraging feedback.
Toward sustainable and ambient support
During our user study, we observed several positive spillover effects of INA on participants’ daily lives, highlighting their potential significance for everyday technology use. However, given the short one-week deployment, the long-term effects of INA remain unstudied. Future work could extend the deployment period and incorporate systematic follow-up studies to assess the application’s generalizability to broad populations, durability, and broader impacts.
Finally, we envision a future version of INA as a form of ambient assistance, where the support is seamlessly embedded in the user’s environment. By developing a system that can be applied across different devices and building a cross-device platform, INA might evolve into a seamless assistant that provides support implicitly, demanding less explicit input, and proactively acting to collaboratively help the user fulfill their intended task (e.g., by hiding irrelevant windows at the start of a session or opening an application it is confident will be needed). Such developments would enable the system to facilitate more intentional and sustainable well-being in everyday life.
9. Conclusion
We present a novel AI-powered assistant system, the Intent Assistant (INA), designed to help users carry out their intended activities in digital environments. We evaluated its ability to detect distraction through a technical evaluation on a custom dataset and validated its effectiveness in a three-week in-the-wild deployment, followed by extensive analysis. Our findings show that INA reduces time spent on intention-irrelevant activities and improves focused immersion compared to simple reminder and logging only applications. User experience analyses further revealed that real-time, context-aware notifications and explicit intention input functioned as core mechanisms fostering reflection and self-regulation. Participants also reported supportive and encouraging experiences. We believe that INA demonstrates the potential of proactive yet autonomy-preserving AI interactions to foster intentional and mindful digital living, paving the way for deeper societal connections between humans and AI.
Acknowledgements.
This material is based upon work supported by the Google Cloud Research Credits program. This work has taken place in part in the Rewarding Lab at UT Austin. During this project, the Rewarding Lab has been supported by NSF (IIS-2402650), ONR (N00014-22-1-2204), ARO (W911NF-25-1-0254), Emerson, EA Ventures, UT Austin’s Good Systems grand challenge, and Open Philanthropy.
Appendix:
State Your Intention to Steer Your Attention: An AI Assistant for Intentional Digital Living
Appendix A System Design Details
We provide details regarding the system designs. The descriptions include the full prompt we used for the system and more details regarding the implementations. We note that we get aid from LLMs on designing prompts, such as fixing grammar or organizing the format so that the LLMs can better understand.
A.1. Clarification
INA incorporates a process of clarification that is conducted to reduce the ambiguity in user-stated intentions for each session. In detail, the process is two-fold: Q&A and formalizing the input text for the prompt. During Q&A, LLMs ask questions to the user to clarify the user’s goal and plan. Then, the answers to the questions are processed, resulting in 10 possible activities predicted by the LLM. These 10 possible activities are included in the input prompt for guiding distraction detection of LLMs as the activity unfolds, serving as candidate reference points for interpreting and disambiguating the user’s ongoing activity, and remain unchanged throughout each session. We provide the full prompts used for clarification below.
Prompt for clarification questions
The text below describes the prompts used for generating clarification questions.
[…]
You are a helpful assistant that engages in a multi-turn conversation to better understand the user’s intention.
[Input] The user stated: “{stated_intention}”
[Goal] Ask one clarification question to make the intention specific and actionable. Questions may concern the specifications of the target item (if shopping), a specific location (if planning a tour), or the specifications of tools the user plans to use (e.g., Slack, e-commerce websites). When asking, guide the user with concrete examples so they can understand and answer easily. Cover diverse aspects of the intention (targets, tools, and related subtasks the user may perform).
[Rules]
1) The objective is to obtain information about the user’s plan or activities related to the stated intention.
2) Focus on questions the user can answer easily about actionable activities (e.g., what tools they will use).
3) Prefer breadth over follow-ups: ask about diverse subtasks or related jobs (e.g., collaborating with peers → communication tools; collecting resources by web search → browsers).
4) Avoid abstract aspects (liking, beliefs, interests); stick to clear points (purpose, plan, activities).
5) Ask for details of the activities (e.g., what kind of essay, what kind of video, which physics topics, what information for “preparing travel to Paris”).
6) Avoid yes/no questions; require brief specific details (e.g., instead of “Will you buy online?”, ask “Which online websites do you plan to use?”).
[Context] Previous Q&A (if any):
First_QA: {first_question_and_answer}
Second_QA: {second_question_and_answer}
[…]
Prompt for processing answers
The text below describes the prompts used for processing answers to the questions and generating 10 possible predictions of user activities.
[…]
System Prompt
You are an assistant that expands simple activity descriptions into diverse, concrete alternatives.
[Guideline] Given a simple activity (e.g., ”Find a jacket for men”), generate 10 variations of the activity description.
[Rules] Use the clarification Q&A with the user. Output must be valid JSON with keys ”1” through ”10”. Keep each sentence concise (at most 9 words). Keep variations non-redundant. Avoid introducing brands/apps/sites unless supported by the clarifications.
[Output structure] Variations 1–3: broader or generalized expressions. Variations 4–6: slightly more specific or rephrased versions using only the original information. Variations 7–10: realistic user actions likely performed when carrying out the activity (e.g., read reviews, search on shopping apps).
[Clarification questions] The following block summarizes clarification Q&A. Use it to augment the intention; if the user specified a website/tool, you may include it (especially in 7–10). {clarification_block}
[Example] Activity: ”Find a jacket for men”
”1”: ”Shopping”,
”2”: ”Online shopping”,
”3”: ”Browse for clothing”,
”4”: ”Search for jackets”,
”5”: ”Look up men’s jackets in an online store”,
”6”: ”Navigate a shopping app to find a jacket”,
”7”: ”Use a search engine to locate jackets”,
”8”: ”Read customer reviews on shopping sites”,
”9”: ”Compare jacket prices across online stores”,
”10”: ”Watch jacket review videos on YouTube”
[Input] Activity: “{stated_intention}”
[…]
A.2. Distraction Detection
INA detects whether the users are on-task or off-task during their activities, based on their activity logs consisting of screenshots and metadata of the application in use. To holistically examine the activity, we use the common-sense reasoning of LLMs and prompt them to consider various aspects of the given information. Also, we prompt the LLM to follow certain output format, eliciting their reasoning capability. For stability, the output scores are discretized into 0.2 increments, ranging from 0.0 (perfectly relevant) to 1.0 (completely irrelevant). We provide the full prompt used for distraction detection below.
Prompt for distraction detection
The text below describes the prompts used for detecting distraction.
[…]
System Prompt
[General Instruction] You are a friendly AI coach with balanced sensitivity to task focus and a neutral communication style. The user’s current intention is provided as [intention: {task_name}]. Help users stay mindful of their task while providing feedback that matches your assigned tone and sensitivity. Consider the specific nature of their task when giving suggestions and feedback. For example, given a task of shopping, the user may watch reviews of several items; or, given a task of writing a report, the user may discuss with peers.
[Clarification Context] Given [intention: {task_name}] and the clarification Q&A, the following augmented-intention items describe plausible activities the user may perform. Use this context for more accurate classification. {list_of_expansion_intention}
[Key Instructions for Evaluating Relevance]
1) Predict Intent: first infer the likely intent behind the current behavior from the provided information.
2) Examine Details: attend to specific cues (e.g., conversation context, YouTube video title).
3) Analyze Beyond Keywords: do not judge by surface category (chat/video/email); decide whether it serves the task.
4) Bridge Indirect Relevance: treat searching, communicating, or researching as relevant when they support the task.
5) Be Certain for Scores: use extreme scores only with clear evidence; otherwise choose an intermediate value.
[Scoring Guidelines]
0.0 — Perfectly relevant: clearly aligned with the task (e.g., writing a report; coding for a project; purchasing a specific item).
0.2 — Mostly relevant: indirect but necessary (e.g., watching a tutorial; reading a related article; discussing with peers).
0.4 — Somewhat relevant: helpful but not essential (e.g., watching a review or a loosely related discussion without clear context).
0.6 — Somewhat irrelevant: unclear whether it supports the task (e.g., browser start page; mixed YouTube thumbnails; finding a file in Finder).
0.8 — Mostly irrelevant: little to do with the task, minimal incidental benefit (e.g., casual video unrelated to the goal; off-topic chat).
1.0 — Completely irrelevant: clear distraction/off-topic (e.g., gaming during study; social media during work; random entertainment browsing).
[Message Instruction]
Always write the user-facing message in polite. If multiple programs are visible, assume the frontmost app reflects what the user is primarily viewing. When the screen provides little information, ground the message in the frontmost app name or URL.
[Message Writing Guidelines]
Use a warm, supportive tone. State what the user is doing simply and clearly. When on-task, share the observation and praise: “I can see you’re focused on {details}.” “Nice progress—this directly supports {details}.” “Keep going with {details}.” When off-task, nudge gently and offer a concrete suggestion: “It looks like {details} may be pulling you away.” “Your focus seems to be drifting to {details}.” “How about switching back to {detailed_suggestion}?” “Let’s refocus on {detailed_suggestion}.” Keep messages short, positive, and momentum-preserving.
[IMPORTANT Rules]
Return only the JSON object.
[CURRENT SCREEN CONTEXT]
Currently active application: {application_name}.
Current URL: {url}.
Please analyze the screenshot considering this context information.
[…]
A.3. Message Generation
INA provides notifications to the users to encourage or discourage the user activities depending on the context. These messages are designed to be both informative and motivating, reflecting the current activity of the user in a supportive tone.
Prompt for message generation
The text below describes the prompts used for the message generation.
[…]
[Message Instruction] While generating the notification message, please follow these guidelines:
- Act as a friendly mother figure who is sensitive to the user’s task.
- Specify what the user is doing as concise as possible.
- When the user is focused, try share what you observe and how you think. For example, start with ”I see your focus on details”, ”details will be helpful for details”, etc. Also, try to keep the message positive and encouraging. nice candidates are ”Keep up coding details”, ”Great focus on writing the report details”, etc.
- When the user is distracted, try to warn the user about the distraction. For example, start with ”I detected distraction details”, ”You seem to be distracted by details”, etc. Also, add some suggestions. For instance, add ”Try returning to detailed suggestion”, ”Please focus on detailed suggestions”, etc.
[…]
A.4. Feedback
INA also incorporates a process of feedback that is to reflect the correction from the user according to the notifications from the system. When the user provides correction feedback, the LLM is prompted to reflect on why the user gave such feedback by analyzing the current situation. Based on the reflection, including interpretation of the user feedback, the LLM generates how it can improve the distraction detection for refining the scoring guidelines. Then, the refinement is included in the input prompt for distraction detection so that the LLM refines its scoring mechanism by using this as a context for future prediction. To robustly incorporate the user correction, we also provide rule-based scoring guidelines, which we prompt together with the reflection, so that the LLM consistently adjusts its scoring direction. Concretely, we defined four types of rule-based scoring guidelines (e.g., “Output higher alignment (lower output score)”, “Output lower alignment (higher output score)”, and their converse cases), depending on whether the system’s prior judgment was correct or incorrect. We provide the full prompt used for feedback below.
Prompt for reflection
The text below describes the prompts used for reflection during the feedback process.
[…]
System Prompt
You are a helpful assistant designed to reflect on your previous alignment judgment using the user’s feedback. Your goal is to infer an implicit intention that was not stated but should have been captured, so that the user’s observed activity is explained in a way that aligns with the current task. Analyze the situation where the user reacted to your prior decision. For example, given a stated intention ”Write a research report” and a screen description ”Chatting with a colleague on Slack,” you may have classified it as a distraction with the rationale ”This appears to be casual conversation, not task-related.” If the user then indicates dislike of your judgment, you should reflect and output an implicit intention such as ”Discuss with a colleague to gather sources for the report,” which aligns the activity with the task.
[Stated Intention] {stated_intention}
[Your Response] A low score indicates you judged the activity aligned with the intention. {assistant_response}
[User Feedback] {user_feedback}
[Task] Reflect on why the user might have expressed this feedback. Consider what implicit intention or subtle task-related reasoning the user might have had that you did not consider. Then propose a policy-adjustment strategy to better align future judgments with the user’s task. The policy statement should follow this format: “Output high/low alignment (low/high score output) for [specific activity with detailed contents] when detected.”
[Output] Return only a JSON object with the following keys:
- ”analysis_assistant_response”: whether your previous response reflected high alignment (low score) or low alignment (high score).
- ”user_activity_description”: a short noun-phrase describing the on-screen activity (e.g., ”YouTube homepage with diverse video thumbnails”, within 20 words).
- ”analysis_user_feedback”: two short sentences (each $\leq$ 10 words) explaining what/why the user liked or disliked your judgment.
- ”user_implicit_intention_prediction”: a short sentence ($\leq$ 10 words) predicting an implicit intention, starting with a verb (e.g., ”Watch review before purchase”).
- ”policy_adjustment”: one sentence following the required policy format above.
[…]
A.5. System Architecture and Implementation
The INA system follows a client-server architecture, consisting of a client application running on the user’s local machine and a backend hosted on a cloud platform. We note that the implementation was aided by the LLM, especially when we code in Python.
Client Application
The client is a Python-based desktop application that is responsible for data transmission to and reception from the back-end.
  • Data Transmission: Transmits user data required for server-side analysis. This data includes (a) the user’s intention and related chat history, (b) periodically captured screen images, and (c) computer usage logs.
  • Interaction Processing: Receives the analysis results of the server, displays them in the UI, and transmits the user feedback back to the server. Key UI functions include presenting questions for intention clarification, displaying notifications after misalignment, and receiving user feedback.
Backend Infrastructure
The backend is developed in Python and deployed on a serverless platform (Cloud Run), which scales according to traffic. Its primary function is to synthesize data from the client and use Google’s Gemini 2.0- Flash model to infer the alignment level between the user’s current activity and their stated intention. It also performs the reflection process required for intention clarification and feedback processing via the LLM.
Data Infrastructure
Data is stored based on its role as follows:
  • Firestore: Stores structured data such as anonymized IDs, intentions, and feedback.
  • Cloud Storage: Stores unstructured data, such as de-identified screen images.
  • BigQuery: Utilized as a data warehouse for long-term pattern analysis.
Appendix B Evaluation of Distraction Detection Details
In this section, we provide details related to the evaluation of distraction detection. The details include more explanation for building IntentionBench and experiments with real-world usage data.
B.1. IntentionBench Details
IntentionBench simulates the realistic user workflows, including the transitions between on-task and off-task activities by synthesizing mixed sessions based on focused sessions. The focused session refers to a record of an activity, where the collector acts as a user and sticks to the given instruction without being distracted. On the other hand, a mixed session includes both activities that are related to a given intention and unrelated to it. The detailed process of generating mixed sessions and collecting focused sessions is as follows.
The focused sessions are collected by two collectors (i.e., the authors). Each focused session was generated by executing 50 distinct task instructions, listed in Table 3 (the instruction column). These instructions span a diverse set of real-world activities, systematically covering web shopping, tour planning, studying, working, and entertaining. Importantly, the collectors are instructed to exploit diverse applications and websites within a single activity, while still adhering to the given instructions. For example, given an instruction “Find a local restaurant”, the collector performed (1) discussing options with a partner on WhatsApp, (2) searching nearby places on Google Maps, and (3) checking reviews on Yelp. In this way, a single task instruction naturally expands into heterogeneous user behaviors across platforms. Then, each collected focused session was divided into smaller segments, with boundaries defined by natural transitions such as switching applications or navigating to new websites (e.g., 3 segments with the “Find a local restaurant” activity record). Also, before executing the instruction, the collectors performed the clarification process, answering two clarification questions generated by the LLMs. These questions often ask what types of applications or websites the user will employ, or what specific items the collector is interested in. The result of this Q&As, including the 10 possible activities that LLM predicted (see Appendix A.1), is also prepared and included in the input prompt during evaluation (especially, for the conditions including clarification feature).
Each mixed session is synthesized by mixing two focused sessions. Specifically, we sample two focused sessions and randomly reorder their segments. Then we label each mixed session with instruction from the first focused session serves as the user intention: segments from this focused session were labeled on-task, while segments from the second were labeled off-task. For the user intention label, we used a relabeled version of each instruction described in Table 3, rather than the instructions used for focused session collection, to reflect the abstract and ambiguous nature that typical users would prefer (as revealed in our formative study). While creating the IntentionBench, we did not include the application metadata, as we considered the screenshot and user intention as primary signals for distraction detection. Also, to reflect configurations that we adopt during deployment, we downsampled the collected records so that each data point accounts for two seconds.
# Instruction Relabeled version
1 Order kids’ books online Buy books
2 Find an affordable TV online Buy a TV
3 Order skincare products Order cosmetics
4 Shop for a BBQ party Prepare for a party
5 Find a jacket for men Buy clothes
6 Find a workout gear Look for workout clothes
7 Buy kitchen appliances Order kitchenware
8 Purchase office supplies Order office supplies
9 Buy home cleaning products Buy daily necessities
10 Order pet supplies Order dog snacks
11 Plan a winter trip abroad Plan a winter trip
12 Plan a summer trip abroad Plan a summer vacation
13 Plan a camping Prepare for camping
14 Find a local restaurant Search for restaurants
15 Find local hiking spots Prepare for hiking
16 Plan a department store shopping Make a shopping plan
17 Find beach destinations Plan a beach trip
18 Plan a dessert tour Go on a dessert café tour
19 Look for amusement parks Do ticketing
20 Schedule a city walking tour Plan a walking tour
21 Coding practice Practice coding
22 Explore ancient Greek myth Study Greek myth
23 Go over biology topics Study biology
24 Read economics news articles Read economic news
25 Learn about quantum mechanics Study quantum mechanics
26 Read history articles on Roman empire Study Roman history
27 Read about modern architecture Study modern architecture
28 Practice a foreign language Study a foreign language
29 Study semiconductor concepts Study semiconductors
30 Derive diffusion model objectives Study diffusion models
31 Survey research papers Read papers
32 Write a business report in a document Write documents
33 E-mail task Handle emails
34 Developing a software Do development (coding)
35 Budget planning on business trip Prepare for a business trip
36 Track team productivity Manage a project
37 Create slides for project proposal Make a PowerPoint presentation
38 Prepare lecture slides Prepare lecture materials
39 Create a company profile slides Create company introduction materials
40 Build a SaaS webpage for companies Develop a website
41 Chill by watching YouTube shorts Watch YouTube
42 Watch stand-up comedy Watch fun videos
43 Play a strategy game Play games
44 Read a fiction novel Read novels
45 Explore new music albums Listen to music
46 Watch a sports event Watch soccer videos
47 Watch movie previews Watch a movie
48 Find a trending Netflix series Watch Netflix
49 Browse social media trends Surf the web
50 Watch an e-sports highlight match Watch e-sports
Table 3. Instructions of 50 tasks and relabeled versions of them. The contents in the instruction column were used for collecting focused sessions, while the contents in the relabeled version were used for labeling mixed sessions.
B.2. Evaluation Details
We evaluate Gemini 2.0-Flash with a temperature of 0.1. The top_p value is set to be 1.0, and the top_k value is set to be 32. We set the max_output_tokens to 512. The other parameters are fixed to be the default. Given a mixed session, the input for the model is the user’s intention and the screenshot image, without application metadata (see Appendix B.1), and the output is a distraction score. Specifically, the model is prompted to output scores that are discretized into 0.2 increments, ranging from 0.0 to 1.0. Then, the final prediction is determined by applying a threshold of 0.5, where scores below 0.5 are treated as on-task and scores of 0.5 or above are treated as off-task. We use the same prompt described in Appendix A.
For evaluating different conditions, we selectively include the interactive features of INA: clarification and feedback. When employing the clarification feature, the result of Q&As corresponding to the user’s intention, i.e., the instruction used for collecting the focused session, is included in the prompt. When incorporating the feedback feature, the user correction is simulated by correcting every false positive. Concretely, whenever the system predicts a distraction but the ground-truth label indicates that the user was actually on-task (i.e., a false positive), this is treated as if the user had expressed ‘incorrect’. In such cases, the system triggers the LLM for refinement (see Appendix A.4). This step asks the model to analyze why its earlier judgment was misaligned with the user’s actual intention, to infer a plausible implicit intention that better explains the activity, and to propose a simple policy adjustment so that similar mistakes are avoided in the future. The outputs of this reflection are fed back into subsequent prompts as additional context, guiding the model to make more accurate predictions in later steps of the same session. Since the real user would not correct all the false positives the model predicts, this heuristic forms an upper bound.
Figure 9. User interfaces for INA (top), simple reminder (middle), and logging only (bottom); INA and simple reminder conduct an Intention alignment rating survey at the end of the session.:
B.3. Evaluation with Real-World Usage Dataset
To validate the detection distraction capability of INA in practice, we perform an additional study by creating a real-world usage dataset. To create the real-world usage dataset, we adopt real user session records from participants for the user study (detailed in Section 6.1). Given the pool of activity session records collected during deployment, we randomly sample 60 data points for each participant, resulting in a total of 1,320 data points (approximately 0.73 hours). Each data point comprises the current screenshot, application metadata, and corresponding user-stated intention collected every 2 seconds, but whether the user was on-task or off-task at each timestep remains unclear. Therefore, three labelers (i.e., the authors) independently labeled whether the user is on-task or off-task, and the ground-truth labels were finalized by majority voting.
With a real dataset, we compared the prediction of INA collected during deployment with the prepared ground-truth label. Here, the evaluated system incorporates all the design components. Performance was measured using the accuracy and balanced accuracy, computed as the average of the true positive rate and the true negative rate. This is because the real data exhibits an imbalanced class distribution. In particular, the data contained substantially more on-task data, as users were already benefiting from the support of INA during deployment.
Then, we study the efficacy of INA in detecting distraction in practice using the real dataset. We observed INA achieves an accuracy of 0.899 in the real dataset, similar to that in IntentionBench with 0.878. Also, INA achieved a balanced accuracy of 0.815 in the real dataset, while that in IntentionBench is 0.865. This suggests that the system attains practical viability even under the noisy and ambiguous conditions of real-world usage. We further build upon this finding as a foundation for subsequent analyses of user behavior and system impact, mainly discussed in Section 7.
Appendix C User Study Design Details
In this section, we provide details on the user study. The information includes an explanation of pre-survey, baselines, survey and interview instruments, and user experience survey questionnaires.
C.1. Pre-survey
For the user study, we conducted a pre-survey to collect participant demographics, assess their computer usage habits, and measure self-regulation.
Demographics and Logistics.
Participants provided their name, phone number, email, age, gender, occupation, and computer specifications. They also reported their average daily computer usage and confirmed their availability for the study. Finally, they provided informed consent.
Usage Habits and Perceptions.
The following questions were rated on a 5-point Likert scale or as open-ended/multiple-selection responses.
  1. (1)
    Technology Proficiency: I quickly become familiar with and can comfortably use new technologies or applications.
  2. (2)
    AI Agent Affinity: I feel comfortable using AI agents (e.g., chatbots, voice assistants) and do not feel resistant to them.
  3. (3)
    Primary Computer Use: For what purposes do you primarily use your computer?
  4. (4)
    Frequency of Getting Sidetracked: How frequently do you find yourself engaging in activities different from your original intention?
  5. (5)
    Improvement Goals: Which aspects of your current computer habits would you most like to improve?
  6. (6)
    Common Distractions: When you get sidetracked, what activities do you typically engage in?
  7. (7)
    Problematic Habits: Please describe any digital device habits you consider ”problematic.”
  8. (8)
    Reasons for Distraction: What is the primary reason for getting sidetracked during work or study?
  9. (9)
    Expectations for AI Agent: Please describe any expectations or concerns about using an AI agent to manage digital usage.
Self-Regulation Scale.
Participants rated their agreement with the following statements on a 5-point Likert scale (1=Strongly Disagree to 5=Strongly Agree).
  1. (1)
    I usually monitor my progress toward my goals.
  2. (2)
    I have a hard time making decisions.
  3. (3)
    I get distracted easily, even when I have a plan.
  4. (4)
    I often realize the consequences of my actions too late.
  5. (5)
    I believe I can achieve the goals I set for myself.
  6. (6)
    I tend to procrastinate on decisions.
  7. (7)
    I often fail to notice when I’ve had enough of something (e.g., alcohol, food, sweets).
  8. (8)
    I am confident that I can make a change if I set my mind to it.
  9. (9)
    When I decide to make a change, I sometimes feel overwhelmed by too many options.
  10. (10)
    Even when I decide to do something, I find it difficult to see it through to the end.
  11. (11)
    I seem to repeat the same mistakes frequently.
  12. (12)
    If I have a plan that is working well, I can stick to it.
  13. (13)
    When I make a mistake, I learn from the experience.
  14. (14)
    I try to set my own standards (personal principles) and live by them.
  15. (15)
    When I see a problem or difficulty, I try to find a possible solution quickly.
  16. (16)
    I find it difficult to set goals that are right for me.
  17. (17)
    I consider myself to have strong willpower.
  18. (18)
    When I try to change something, I carefully check my progress.
  19. (19)
    It is not easy for me to make a concrete plan to achieve a goal.
  20. (20)
    I am not easily swayed by temptation.
  21. (21)
    I set goals and check how close I am to achieving them.
  22. (22)
    I tend not to pay close attention to what I am doing.
  23. (23)
    Even if my current method is not working, I tend to stick with it instead of changing it.
  24. (24)
    When I want to change something, I tend to look for several alternatives.
  25. (25)
    When I have a goal, I can create a plan to achieve it.
  26. (26)
    Once I decide to change something, I consistently monitor the process.
  27. (27)
    I am often unaware of what I am doing until someone else points it out.
  28. (28)
    I usually think carefully before I act.
  29. (29)
    I learn from my mistakes.
  30. (30)
    I have a clear idea of the person I want to become.
  31. (31)
    Before making a decision, I carefully consider the consequences of each choice.
C.2. Baselines Details
We describe the interfaces of the baseline systems in Figure 9. In the simple reminder, users also begin by typing their intention into a text field and pressing the ‘Start’ button. Unlike INA, no clarification process takes place, and the session begins immediately. During the session, the system provides periodic reminders by displaying a static notification stating the user’s stated intention (e.g., “Your intention is Study”). At the end of the session, triggered by the user pressing the ‘Stop’ button, users are asked to complete an intention alignment rating survey.
In the logging only, users press ‘Start’ to begin a session without inputting their intention. No clarifications or notifications are provided during the session. The system records user activities until the user presses the ‘Stop’ button. This baseline serves as a minimal setup to isolate the effect of recording intentions without additional reminders or feedback.
C.3. Survey and Interview Instruments
Scope Statement / prompt (English translation of Korean wording)
User experience questions  (Jörke et al., 2025)
All applications The program understood my intentions. (modified GPTCoach Q9)
All applications The program helped me act on the activities I intended to do. (modified GPTCoach Q13)
All applications I felt supported by the program. (modified GPTCoach Q5)
All applications The program made me feel capable of controlling my digital activities. (modified GPTCoach Q6)
All applications The program used my data in a meaningful way. (modified GPTCoach Q12)
All applications I worried my personal information might leak.
All applications The program gave me new insights into my computer use. (modified GPTCoach Q16)
All applications Using this program interfered with my workflow. (rev.)
INA +simple reminder Notifications helped me act according to my intentions.
INA +simple reminder Notifications motivated intended activities.
INA +simple reminder Notifications arrived at appropriate times.
INA +simple reminder The program tailored its messages to my intentions and activities.
INA +simple reminder The program provided clear explanations for its message.
INA +simple reminder Notifications felt burdensome. (rev.)
INA only The program adapted its behaviour based on my feedback.
Focused-Immersion 5-item scale (Agarwal and Karahanna, 2000)
All applications (FI1) While using the computer I am able to block out most other distractions.
All applications (FI2) While using the computer , I am absorbed in what I am doing.
All applications (FI3) While on the computer , I am immersed in the task I am performing.
All applications (FI4) When on the computer , I get distracted by other attentions very easily. (rev.)
All applications (FI5) While on the computer , my attention does not get diverted very easily.
Open-ended Survey
All applications Was there a moment when the program helped you focus or carry out tasks as intended? If so, please explain.
All applications Has the program helped you focus or carry out tasks as intended? If so, please describe with examples (context and how the program helped).
All applications What did you like or dislike about the program?
All applications What insights, if any, have you gained about your digital habits or computer use?
10-20 minute semi-structured video interview
All applications What was your overall experience using the program?
All applications Usage context: When/where did you open [App] most, and what were you trying to do?
All applications Did [App] help you live as intentionally as planned? If not, why?
All applications Anything about [App] that pleasantly or worryingly surprised you?
All applications Would you keep using [App] One decisive reason to keep One thing to fix first
INA +simple reminder Impact of notifications: one helpful moment and one ill-timed or burdensome moment.
INA only Intent understanding — one “hit” and one “miss”.
INA only Effect on personal agency — did the agent enhance or undermine it? Example.
INA only Perceived adaptivity — one case where adaptation worked (or failed).
Table 4. Full list of subjective items used in the weekly survey (Likert 1–5) and the post-week phone interview. In brackets we note the original source question that inspired each item.
We provide the questionnaires used in the survey and interview in Table 4. The survey and interviews aim to understand how three systems influenced users’ intentional digital activity and what their experiences of usage were. They ask whether the system understood intentions and supported focused activity, and whether notifications were perceived as helpful, timely, motivating, or burdensome. The questions also ask about the effects on users’ sense of control, workflow, and immersion, and concerns, such as privacy or unexpected disruptions. Finally, the interviews explore broader experiences, including insights gained about the users’ digital habits and perceptions of adaptivity or autonomy. In our analyses, we focused on the interview responses and excluded those from the open-ended survey, as their contents largely overlapped.
Appendix D Further Analyses
In this section, we present further analyses, auxiliary to the findings in our study.
D.1. Proportion of Off-task Time for Categorized Intention
We categorized user-stated intentions into productive, non-productive, and uncategorized categories by directly mapping the ICATUS 2016 major divisions (United Nations Statistics Division) (United Nations Statistics Division, 2021). As shown in Table 5, Divisions 1–6 were defined as productive activities and Divisions 7–9 as non-productive activities. Ambiguous or mixed cases that did not clearly fall into either category were assigned to an uncategorized class. Two independent researchers coded all intentions. To assess inter-rater reliability before reconciliation, we computed Cohen’s $\kappa$, which indicated almost perfect agreement ($\kappa=0.9222$). Remaining disagreements were resolved through discussion. Figure 10 shows the mean LLM-estimated off-task ratios for productive and non-productive intentions.
Category ICATUS Major Division
Productive 1. Employment and related activities
2. Production of goods for own final use
3. Unpaid domestic services for household and family members
4. Unpaid caregiving services for household and family members
5. Unpaid volunteer, trainee and other unpaid work
6. Learning
Non-productive 7. Socializing, communication, community participation, religious practice
8. Culture, leisure, mass-media and sports practices
9. Self-care and maintenance
Uncategorized Ambiguous or mixed cases
Table 5. Mapping of ICATUS 2016 major divisions to productive vs. non-productive categories.
(a) Productive intentions
(b) Non-productive intentions
Figure 10. Mean LLM-estimated off-task ratio ($\downarrow$) for productive intentions and non-productive intentions. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are more favorable.:
D.2. Relation between Intention Alignment Rating and LLM-estimated off-task ratio
We examined the relation between intention alignment rating and the LLM-estimated off-task ratio, and as shown in Figure 11, higher intention alignment ratings are associated with lower off-task ratios.
Figure 11. Mean LLM-estimated off-task ratio shown by intention alignment rating (1–5). Error bars denote 95% confidence intervals across participants (user-averaged).:
D.3. User Experience Survey Results
We analyzed the user experience survey by category, dividing items as shown in Table 6, where the specific questions for each category can be found. User responses to individual items are presented in Figure 12. Items marked with (rev.) are reverse-coded, so that higher values across all items indicate more positive meanings. INA outperformed both simple reminder and logging only across all dimensions. Participants reported that INA better understood their intentions, supported them in carrying out intended activities, and delivered notification messages of consistently high quality. However, in Q8 and Q12, some participants also experienced these supportive functions as distracting or burdensome.
Factor ID Full question sentence Cronbach’s $\alpha$
Support Q1 The program understood my intentions. 0.81
Q2 The program helped me act on the activities I intended to do.
Q3 I felt supported by the program.
Q4 The program made me feel capable of controlling my digital activities.
Q5 The program used my data in a meaningful way.
Q6 I worried my personal information might leak. (rev.)
Q7 The program gave me new insights into my computer use.
Workflow Disruption Q8 Using this program interfered with my workflow. (rev.)
Message Effectiveness Q9 Notifications helped me act according to my intentions. 0.78
Q10 Notifications motivated intended activities.
Q11 Notifications arrived at appropriate times.
Q12 The program tailored its messages to my intentions and activities.
Q13 The program provided clear explanations for its message.
Table 6. Survey items grouped by factor. The categorization into Support, Workflow Disruption, and Message Effectiveness was validated by internal consistency, expressed with Cronbach’s $\alpha$. Values above 0.70 indicate acceptable reliability, supporting the appropriateness of the factor structure. Cronbach’s $\alpha$ was not computed for groups with a single item. Reverse-coded items are marked with (rev.).
Figure 12. Results of 15-item user experience survey responses by application on a 5-point Likert scale. Means and standard deviations are shown on the right; (rev.) indicates reverse-coded items. ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively.:
D.4. Interview Coding Reliability Thematic Analysis
A total of 66 interview recordings were transcribed into text, as 22 participants each took part in interviews about three applications. We conducted a codebook thematic analysis that began with inductive coding and iterative refinement. Two coders, including the first author, independently generated initial codes while reading the transcripts and reached consensus on a preliminary codebook. Using this codebook, the coders independently coded the data, compared results, and refined definitions and boundary rules by merging and specifying codes based on observed response patterns.
To assess intercoder reliability efficiently, the coders independently coded nine transcripts randomly sampled in equal proportions from each application. For each coder, we constructed a binary presence/absence matrix and computed Cohen’s $\kappa$, which indicated substantial agreement ($\kappa=0.77$).
Afterward, the two coders each coded the remaining transcripts and cross-checked the results, and finally, the first author reviewed and explained the outcomes to the team. The final codebook comprises 24 codes, each with positive, neutral, and negative subcategories, and is summarized in Table 2.
Coding principles were threefold. First, codes were assigned based on the semantic surface of utterances to minimize interpretation. Second, multiple codes could be applied to the same unit. Third, for descriptive statistics, we used respondent-level binary aggregation, such that repeated mentions of the same code by the same participant were counted once.
D.5. Open-ended Survey Coding Descriptive Analysis
Theme Code Definition logging only (%) simple reminder (%) INA (%)
Effects Return-to-task Notifications interrupt off-task behavior and redirect to intended activity 0.0 23.0 64.0
Induced immersion Intention setting ($\rightarrow$ Q&A interaction) $\rightarrow$ self-evaluation induces deeper immersion 0.0 27.0 32.0
Deterrence via perceived observation Perceived observation (system or staff) deters off-task behavior 14.0 9.0 9.0
Visual salience System presence / color cues enhance attentional focus 27.0 32.0 9.0
No effect No benefit when features feel insufficient/misclassified or hinder use 59.0 14.0 9.0
Pros Efficient time use Reduced unnecessary use; more purposeful time allocation 9.0 14.0 14.0
Intention declaration effect Entering/refining intentions organizes plans; deepens immersion 0.0 18.0 18.0
Real-time notifications Context-appropriate prompts aid return from off-task; convey support 0.0 0.0 36.0
Activity logging & visualization Reviewing what was done and for how long is helpful 0.0 14.0 14.0
Non-intrusiveness No obtrusive features; does not interfere with tasks 23.0 0.0 0.0
Cons No observable benefit Little/no help; features feel insufficient 68.0 5.0 0.0
Excessive notifications Prompts too frequent/nagging; disrupt activity 0.0 9.0 27.0
Insufficient notifications Prompts absent or too infrequent; limits effectiveness 5.0 23.0 0.0
Misclassified notifications Inaccurate inference leads to context-incongruent messages 0.0 0.0 27.0
Reliability & safety concerns Crashes, lost intentions, privacy concerns 14.0 27.0 27.0
UX degradation Impaired workflow (e.g., dual monitors blocked, partial occlusion) 0.0 27.0 23.0
Insights Frequent social-media checking Frequent checking of messengers/SNS 14.0 9.0 23.0
Low sustained attention Short attention spans; drift within $\sim$20 minutes 27.0 36.0 36.0
High/improved attention Rarely straying off-task; reduced off-task proportion 9.0 18.0 18.0
Notification sensitivity Stress / heightened reactivity when prompts appear 0.0 5.0 5.0
Frequent video consumption Frequent YouTube/OTT video use 0.0 5.0 5.0
Routine awareness Awareness of app-specific usage amounts/sequences 0.0 0.0 9.0
None/unsure No clear realizations or unsure 50.0 41.0 0.0
Table 7. Open-ended survey themes with percentages by condition (logging only/ simple reminder/ INA). Values are the proportion of participantsN=22N=22 reporting each code.
We applied descriptive coding to analyze the open-ended survey responses. This method assigns short labels to each response segment to summarize the experiences or perceptions reported by participants. The final codebook and results are presented in Table 7. The results illustrate that logging only provided unobtrusive logging with little support, simple reminder added reminders with mixed sufficiency, and INA delivered the strongest refocusing effects while introducing new challenges in precision and reliability.
Specifically, INA frequently reported of “Return-to-task” (64.0% vs. 23.0% simple reminder, 0.0% logging only), showing that adaptive prompts often helped participants refocus. Mentions of “Induced immersion” (32.0% vs. 27.0% simple reminder, 0.0% logging only) further indicate that intention setting and clarification deepened immersion compared to logging only. In the positive codes, INA uniquely noted “Real-time notifications” (36.0%), while both INA and simple reminder reported “Intention declaration effect” (18.0%) and similar levels of “Efficient time use” (14.0%) and “Activity logging & visualization” (14.0%).
Furthermore, negative reports diverged by condition. INA showed high rates of “Excessive notifications” (27.0%), “Misclassified notifications” (27.0%), and concerns around “Reliability & safety” (27.0%) and “UX degradation” (23.0%). simple reminder more often noted “Insufficient notifications” (23.0%), while logging only concentrated on “No observable benefit” (68.0%) and the effect-level “No effect” (59.0%). Conversely, logging only participants emphasized “Non-intrusiveness” (23.0%) and “Visual salience” (27.0%), which were rare in INA (0.0% and 9.0%).
Appendix E Privacy and Security Policies
The system was implemented with user privacy and data security as considerations throughout its design. The main policies are as follows.
Data De-identification
Screen images undergo a de-identification process via a Data Loss Prevention service before storage. The original images used for real-time inference are immediately discarded, and only the masked images, with sensitive information such as names and phone numbers removed, are stored.
Access control
Access to data storage is strictly controlled based on the principle of least privilege, using multi-factor authentication and Identity and Access Management policies.
Data Encryption
All data is encrypted in transit between the client and server using the Transport Layer Security protocol. Data at rest is also stored in an encrypted state.